OpenStudy (anonymous):

Is the cube root of 250 rational or irrational? Use 2 different strategies to justify your answer. Please help...

OpenStudy (lgbasallote):

interesting..

OpenStudy (anonymous):

\[\sqrt[3]{250}=\sqrt[3]{125*2}=5*\sqrt[3]{2}\] and the cube root of 2 is irrational so the cube root of 250 is irrational too.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thank you so much, hope it's right!

OpenStudy (anonymous):

Guaranteed. Correct or money back. ;)

OpenStudy (anonymous):

;D